# LocMoE+: Enhanced Router with Token Feature Awareness for Efficient LLM Pre-Training

### v1

@@ A.1 Auxiurary Results
Appendix B Token Feature Distribution
We also validate the feature distribution before and after MoE training shown in Figure 7. We can see before training, all 8192 tokens in one training sample are nearly orthogonal with correlation coefficient near zero, which verifies the isotropy distribution assumption in the first bullet of Remark 3.4. After training, the token features are nearly aligned with correlation coefficien large than $0.8$. We can also observe that neighbouring tokens share similar features, and clear block feature behavior, meaning that the token features are relatively separated and the number of tokens in each cluster is bounded, which somehow matches the distribution assumption in the second bullet of Remark 3.4.
@@ Appendix B Token Feature Distribution
Appendix C Experimental Setup
C.1 Datasets for Training and Fine-Tuning
@@ C.2 Experimental Environment
C.3 Model Configuration
This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and LocMoE+. Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention. It contains 32 sparse expert blocks, with the MoE Feedforward layer comprising 8 experts, and each token selects the top 2 experts for processing. In the application scenarios of this paper, the corpora are generally long texts; thus, the sequence length is adjusted to 32768. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. The batch size is set to 128.
C.4 Evaluation Metrics and Datasets
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval [40] and TeleQnA [41] to evaluate the model’s Chinese language capability.

### v2

@@ Appendix B Missing Proof
Appendix C Token Feature Distribution
We also validate the feature distribution before and after MoE training shown in Figure 8. We can see before training, all 8192 tokens in one training sample are nearly orthogonal with correlation coefficient near zero, which verifies the isotropy distribution assumption in the first bullet of Remark 7. After training, the token features are nearly aligned with correlation coefficien large than $0.8$. We can also observe that neighbouring tokens share similar features, and clear block feature behavior, meaning that the token features are relatively separated and the number of tokens in each cluster is bounded, which somehow matches the distribution assumption in the second bullet of Remark 7.
@@ Appendix C Token Feature Distribution
Appendix D Experimental Setup
Datasets for Training and Fine-Tuning
@@ Appendix D Experimental Setup
Model Configuration
This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and our approach. Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention. It contains 32 sparse expert blocks, with the MoE Feedforward layer comprising 8 experts, and each token selects the top 2 experts for processing. In the application scenarios of this paper, the corpora are generally long texts; thus, the sequence length is adjusted to 32768. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. The batch size is set to 128.
Evaluation Metrics and Datasets
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval (Huang et al. 2024) and TeleQnA (Maatouk et al. 2023) to evaluate the model’s Chinese language capability.